### v1

Abstract
End-to-end learning has shown great potential in autonomous parking, yet the lack of publicly available datasets limits reproducibility and benchmarking. While prior work introduced a visual-based parking model and a pipeline for data generation, training and close-loop test, the dataset itself was not released. To bridge this gap, we create and open-source a high-quality dataset for end-to-end autonomous parking. Using the original model, we achieve an overall success rate of 85.16% with lower average position and orientation errors (0.24 meters and 0.34 degrees). The dataset can be found in https://pan.baidu.com/s/16LSSfF_Qar3e-s_dseDt1g?pwd=wybs.
keywords:
Autonomous Driving, Autonomous Parking, End-to-End Neural Network, Open Dataset, Deep Learning
1 Introduction
The field of autonomous driving has advanced significantly, with a focus on improving autonomous parking. Traditional multi-stage pipelines, including perception, mapping, planning, and control, often suffer from sensor noise, model uncertainties, and rigid handcrafted rules, leading to error accumulation and poor adaptability in complex environments.
End-to-end learning has emerged as a potential alternative by directly mapping sensor inputs to control actions, streamlining the pipeline, reducing reliance on handcrafted features, and improving adaptability. This approach has shown success in lane keeping, object detection, and collision avoidance. Inspired by this, previous work proposed an end-to-end parking method integrating visual inputs and motion states to generate control signals.
However, a key challenge in this domain is the lack of publicly available datasets for training and evaluation. While Yunfan Yang et al.[1] introduced a model and shared a pipeline ranging from data generation to training and closed-loop test, their dataset remains unreleased. To address this gap, we create and open-source an end-to-end autonomous parking dataset based on their environmental setup.
In this paper, we present the iterative process of creating a parking dataset that captures a wide variety of parking scenarios. Using the model architecture proposed by the original authors, we trained the model on the dataset and achieved an overall success rate of 85.16%, which is comparable to the original authors’ result of 91.41%. Additionally, we observed significant reductions in both average position and orientation errors (0.24 meters and 0.34 degrees) compared to the baseline (0.30 meters and 0.87 degrees), indicating improved parking precision.
Our work not only provides a valuable dataset for the research community but also demonstrates the effectiveness of the original model on a brand new and diverse dataset. To encourage further research and development in the field of end-to-end autonomous parking including optimizing algorithm and iterating dataset, we release the latest dataset and comparison experiment results during the dataset iteration process for public access. The main contributions of this paper include:
  • The creation and open-sourcing of a new, high-quality end-to-end autonomous parking dataset based on CARLA, addressing the lack of available data in this domain.
  • The application of the original end-to-end autonomous parking model to this new dataset, achieving competitive results in terms of Target Success Rate, Average Position Error and Average Orientation Error.
  • The release of iteration details for dataset creation, providing a foundation for future research in autonomous parking.
2 Related Works
In this part, the development of Transformer model and BEV model will be reviewed, and based on this, the application of these algorithms in the field of automatic parking will be introduced.
2.1 Transformer model
Transformer models have revolutionized AI since their introduction in the paper “Attention Is All You Need” by Vaswani et al. in 2017 [2]. It was first applied to Natural Language Processing (NLP) where it achieved significant success, such as BERT[3], GPT[4] and T5[5]. In the field of computer vision (CV), Vision Transformers (ViTs)[6] are able to outperform CNNs in image classification tasks by capturing both global and local features and understanding image-wide context. DETR[7] is a successful end-to-end object detection framework with a Transformer encoder-decoder architecture.
Transformer models are also widely applied to autonomous driving. Transfuser[8] and InterFuser[9] use transformers for multi-scale fusion of image and LiDAR features. BEVFusion[10] is a framework that combines camera and LiDAR features in the Bird’s Eye View (BEV) space and uses transformers to fuse multi-modal data, promoting the understanding of the driving environment. UniAD[11] utilizes four transformer decoder-based perception and prediction modules to establish end-to-end autonomy. ParkPredict+[12] uses transformers for waypoint prediction in complicated parking environments. DriveTransformer[13] is a unified transformer framework, introducing task parallelism, sparse representation, and streaming processing to enhance training stability and efficiency. TransParking[14] is a vision-based transformer framework for end-to-end autonomous parking, using expert trajectory training to directly predict future coordinates. E2E Parking[1] sets up a pipeline for data generation, training and close-loop test based on CARLA simulator, implements an LSS-based transformer architecture for imitation learning, predicts control signals directly from RGB images and motion states, and achieves a target success rate of 91%. As an extention, ParkingE2E[15] proposes a camera-based end-to-end parking model that fuses surround-view RGB images in BEV space by target query and predicts trajectory points using a transformer decoder, demonstrating reliability across various scenarios.
2.2 BEV representation
The Bird’s Eye View (BEV) is widely used in autonomous driving for tasks like semantic segmentation and object detection, providing a top-down perspective of the environment. Since 2D image features lack depth, LSS[16] method learns depth distributions and uses cameras’ intrinsic and extrinsic matrices to transform the frustum into BEV. Based on LSS, BEVDepth[17] improves depth estimation with explicit supervision. Transformer models, as in Transfuser[8], are used for BEV generation by fusing LiDAR and image features. ST-P3[18] is an end-to-end vision-based autonomous driving framework that leverages spatial-temporal BEV feature learning for perception, prediction and planning.
2.3 Autonomous driving datasets
The KITTI dataset [19] provides 6 hours of diverse real-world traffic scenarios with synchronized sensor data, including stereo cameras, a 3D laser scanner and GPS/IMU, supporting tasks like stereo vision, optical flow, and object detection. The nuScenes dataset[20], with a full autonomous vehicle sensor suite (6 cameras, 5 radars, 1 LiDAR), provides 1000 fully annotated 20s scenes with 3D bounding boxes for 23 classes, aiming at detection and tracking evaluation with novel metrics. The Waymo Open Dataset[21] consists of 1150 diverse, high-quality 20s scenes with synchronized and calibrated LiDAR and camera data and exhaustive 2D and 3D annotations, used for 2D and 3D detection and tracking tasks. NAVSIM[22] is a benchmarking framework for non-reactive autonomous vehicle simulation, utilizing diverse datasets like nuPlan to evaluate end-to-end driving policies through standardized open-loop metrics, supporting tasks such as planning and trajectory prediction.
For autonomous parking, the PKLot Dataset[23] contains 695,899 images from two parking lots with three camera views under various weather conditions, enabling benchmarking and evaluation of parking vacancy detection algorithms. SUPS[24] is a simulated underground parking dataset featuring multi-sensor data (fisheye cameras, pinhole cameras, LiDAR, IMU, GNSS) with pixel-level semantic labels, applied to tasks such as SLAM, semantic segmentation, parking slot detection, 3D reconstruction and depth estimation. The Dragon Lake Parking (DLP) dataset[12] provides high-resolution aerial video and detailed annotations of vehicles, cyclists, and pedestrians in a large parking lot, supporting research on trajectory analysis, object detection, tracking, and semantic segmentation.
Most of the related datasets are targeted at detection, classification, segmentation and planning, while the mainstream end-to-end autonomous parking methods primarily predict trajectory points. E2E Parking is one of the few studies that directly predict control signals. However, the original authors do not release their dataset, limiting further research on the algorithm, which serves as our motivation for developing a dataset for it.
3 Method
The innovative architecture proposed by the original authors is continued in this work, where a neural network maps inputs to control signal outputs.
3.1 Problem Definition
The system aims to achieve precise vehicle parking control using sensor measurements. The model is trained in a supervised learning framework with expert demonstration dataset $\mathcal{D}=\{\tau^{n}\mid n\in\{1,2,\dots,N\}\}$, where each of the N trajectories $\tau_{n}$ consists of sequential observations and control signals: $\tau^{n}=\left\{\left(\mathcal{X}_{t}^{n},\mathcal{C}_{t}^{n}\right)\right\}_{t=1}^{T^{n}}$. $\mathcal{X}_{t}^{n}$ includes RGB images of four cameras, target slot coordinates and ego vehicle’s motion state in the $n$th trajectory at time step $t$, while $\mathcal{C}_{t}^{n}$ consists of values of throttle, brake, steering and gear state in the $n$th trajectory at time step $t$.
The learning objective is formulated as:
$\arg\min_{\pi}\mathbb{E}_{(\mathcal{X},\mathcal{C})\sim\mathcal{D}}\left[\mathcal{L}\left(\mathcal{C},\pi(\mathcal{X})\right)\right]$ (1)
where $\pi$ is the control policy mapping inputs to control signals, and $\mathcal{L}$ is the loss function.
3.2 Inputs and Outputs
Inputs: The model takes camera images, motion states (velocity and acceleration), and the target slot position. Four surrounding cameras (front, left, right, rear) are positioned to capture the environment. The target slot is initially set by the user and then continuously tracked via segmentation.
Outputs: The model predicts control signals for the next four steps with 0.1s interval. The outputs include normalized acceleration ($\text{acc}_{t}\in[-1,1]$), steering angle ($\text{steer}_{t}\in[-1,1]$), and gear selection ($\text{gear}_{t}\in\{0,1\}$). Inspired by Pix2Seq [25], control prediction is framed as a language modeling task, where signals are tokenized and predicted sequentially. Acceleration and steering are discretized into 201 values $[0,200]$, with 0 indicating full brake/left and 200 indicating full throttle/right. For gear state, 0 stands for forward gear and 200 means reverse gear. The sequence format is:
$S=[\text{BOS},c_{0},c_{1},c_{2},c_{3},\text{EOS}]$ (2)
$c_{n}=[\text{acc}_{n},\text{steer}_{n},\text{gear}_{n}].$
3.3 Network Architecture
The architecture consists of BEV Generation, Feature Fusion and Control Prediction.
BEV Generation: Utilizing the LSS method[16], multi-view images $I\in\mathbb{R}^{3\times H\times W}$ are processed by the backbone EfficientNet-B4 to extract features $F_{\text{img}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ to estimate depth distribution $D$. Feature frustums $F_{\text{frus}}\in\mathbb{R}^{D\times C_{i}\times H_{i}\times W_{i}}$ are then generated, which are projected into a BEV voxel grid and merged to form the final BEV feature map $F_{\text{bev}}\in\mathbb{R}^{C_{b}\times X_{b}\times Y_{b}}$.
Feature Fusion: The BEV feature, target slot, and ego motion are combined into a unified feature. By setting target slot grids as 1 and other grids as 0, the target feature $F_{\text{target}}\in\mathbb{R}^{1\times X_{b}\times Y_{b}}$ is projected into a BEV grid, concatenated with $F_{\text{bev}}$, and processed through ResNet18 convolution layers to obtain $F^{\prime}_{\text{bev}}\in\mathbb{R}^{C^{\prime}_{b}\times L}$, where $C^{\prime}_{b}=C_{b}+1$. Ego motion features $F_{\text{ego}}\in\mathbb{R}^{3\times L}$ are encoded by MLP, representing the information of velocity, x-direction and y-direction acceleration. The fused feature $F_{\text{fuse}}\in\mathbb{R}^{(C^{\prime}_{b}+3)\times L}$ is then fed into a Transformer encoder for enhanced representation.
Control Prediction: A Transformer decoder predicts control signals in an auto-regressive method. The fused feature $F_{\text{fuse}}$ serves as keys $K$ and values $V$ in cross-attention, while the empty sequence serves as queries $Q$. The decoder iteratively generates control tokens, which are detokenized into normalized acceleration, normalized steering angle and gear state, then mapped to control commands.
3.4 Loss
There are two main loss functions and one auxiliary task:
Control Signal Loss: A cross-entropy loss is used to minimize the discrepancy between predicted and ground truth control sequences.
Semantic Segmentation Loss: The model classifies BEV grid objects into vehicle, target slot, and background, supervised with cross-entropy loss.
Auxiliary Task: Inspired by BEVDepth [17], depth prediction is applied to enhance interpretability and improve performance.
4 Dataset Generation
4.1 Implementation details
The simulator is CARLA 0.9.11, which keeps the same as the one in the original paper. The Figure 1 shows the distribution of slots in the parking lot, where the red letter “T” designates the target slot. The red parking slots 2-2, …, 2-16, 3-2, …, 3-16 which end with even number are used for data generation and the brown parking slots 2-1, …, 2-15, 3-1, …, 3-15 which end with odd number are used for closed-loop test. Most of the settings are the same as the original ones. The ego vehicle is initialized at a random position of the path between the second and the third row, heading parallel with the path.
Figure 1: Top view of the parking lot in CARLA
4.2 Problems and solutions
Through preliminary reproduction of dataset and corresponding tests, there exist following problems. The core of dataset optimization is to solve these problems, which is the guideline of subsequent experiments.
4.2.1 Keyboard Control
The original authors used a keyboard for vehicle control with four basic operations: throttle, brake, steer, and gear shift.
  • Throttle: Press W or Up Arrow to increase throttle by 0.05 per frame (up to a maximum of 0.5). Releasing resets to 0.
  • Brake: Press S, Down Arrow, or Space to increase brake intensity by 0.2 per frame (up to a maximum of 1). Releasing resets to 0.
[…]
  • Gear Shift: Press Q to toggle between forward (1) and reverse (-1).
Keyboard control is challenging for parking due to its discrete nature, which does not intuitively correspond to the continuous control inputs required for smooth maneuvering. To improve control, an XBOX controller was implemented to allow direct linear input. Triggers act as accelerator/brake pedals, and joysticks resemble a steering wheel (self-centering when released). This provides smoother, more intuitive control while maintaining the original value ranges and resolutions.
4.2.2 Insufficient scenes and dataset size
Analysis of the provided code indicates that the --random_seed argument in the argparser lacks an explicit int type definition. As a result, the specified random seed value was interpreted as a string, leading to an error when invoked. This suggests that the authors used only the default random seed value (0), with its increment tied to the task index. Consequently, each parking slot was assigned only one single random seed for scene creation, limiting scene variation.
The original authors use random seeds from 0 to 15 for the 16 parking slots. As an improvement, we used random seeds from 0 to 15, assigning each to one parking slot and collecting eight parking routes per slot. The process was then repeated with seeds from 16 to 31 for another 16 slots. Random seeds 32–47 were also assigned to 16 slots, but only four routes were collected per slot.
4.2.3 Lack of parking slot with shadow of streetlight
Our experiments demonstrated that the trained model fails to park the vehicle into a parking slot with streetlight shadows when strictly following the data generation pipeline. This issue arises because only slot 3-9 contain shadows, which are used for model test. During data sampling, the vehicle was never parked into a slot with a shadow, leading to difficulties during testing with the slot 3-9.
To capture parking slots with shadows while adhering to the original authors’ division of training and testing slots, we decided to adjust the solar altitude and azimuth angles. Two scenarios were generated, with the shadow projected onto slots 3-8 and 3-10, respectively. When a shadow is projected onto slot 3-8, random seeds 11 and 27 were assigned to collect 8 routes per seed. Similarly, when a shadow is projected onto slot 3-10, random seeds 12 and 28 were used to collect 8 routes per seed.
4.2.4 Trouble of parking slots near the edge of the parking lot
During testing, when the target parking slots are 2-1 and 3-1 and the vehicle initially faces outward from the parking lot, the ideal maneuver involves driving straight to the edge of the parking lot, then adjusting the steering and reversing into the slot. However, in the training set, parking routes were collected for neighboring slots 2-2 and 3-2, without instances of the vehicle reaching the edge of the parking lot. As a result, during testing, the model encounters previously unseen scenarios when handling slots 2-1 and 3-1, leading to reduced stability.
As a solution, compared to the original speed limits of 12 km/h and 10 km/h for forwarding and backing, the speed is now further limited to 7.5 km/h for both directions. Experiments have proven that this adjustment improves the Target Success Rate.
4.3 Dataset overview
Each route data of the dataset was generated via the script provided by original authors. The sampling frequency is 10 Hz. The data for model input is RGB images from 4 cameras, extrinsic and intrinsic matrix for each camera, target point and motion state of ego vehicle, while control values, depth and segmentation are ground-truth labels. The range and resolution of the control values are listed in Table 1. 5 versions of dataset were created for comparison experiments by one expert driver: Gen 1A, 1B, 1C, 2A, 2B.
Table 1: Range and resolution of control values
Control Range Resolution
Throttle [0, 0.5] 0.1
Brake [0, 1] 0.2
Steering [-0.7, 0.7] 0.1
Reverse {0, 1} /
[…]
In Gen 1, after driving straight, the vehicle stops at a position slightly farther from the target parking slot. It then gradually increases the steering angle while reversing. When both rear wheels cross the extended boundary line of the target slot, the steering angle is increased to its maximum. Subsequently, the steering and throttle values are gradually reduced. As the vehicle’s front approaches alignment, the steering angle is fine-tuned. Once aligned, the throttle is further adjusted in reverse until the vehicle reaches the desired parking position, at which point the brake is fully applied. During the model test for parking slot 2-1 and 3-1, the initialization of the vehicle is outside the parking lot, with half of the body on the road. Therefore, during the data generation for 2-2 and 3-2, original initialization was modified to make the vehicle generated at that position in 10 routes, so that the model would have potential to handle this type of situation.
In Gen 2, the vehicle stops closer to the target parking slot after driving straight compared to Gen 1. It then reverses with the steering angle at its maximum. As the rear approaches the slot, the steering and throttle values are gradually decreased. The subsequent operations are the same as those in Gen 1. There is a slight change for initialization compared to Gen 1, in order to remove the instability caused by roadside shoulder. In Gen 2, the initial positions of parking slots 2-1 and 3-1 during testing are always within the parking lot, whereas for some routes of slots 2-2 and 3-2 during data generation, the vehicle is initialized near the edge of the parking lot to ensure coverage of edge-adjacent scenarios.
The selection of random seeds for each dataset is listed in Table 2. The random seed ranges refer to the seeds employed in the scene generation process, which can change the total number of vehicles and the type of vehicle in each parking slot.
Table 2: Number of trajectories in each dataset with different random seeds
Scenario Trajectories (Train / Val) in Dataset
Gen 1A Gen 1B Gen 1C Gen 2A Gen 2B
Seed 0 to 15 96 / 32 96 / 32 96 / 32 96 / 32 96 / 32
Seed 16 to 31 96 / 32 96 / 32 96 / 32
Seed 32 to 47 32 / 32
Seed 16 to 143 96 / 32  [^1] Gen 2A has more scenes. For example, random seeds 16, 32, 48, 64, 80, 96, 112 and 128 are assigned to the first parking lot for generation.
Slot 3-8 with shadow, seed 11, 27 12 / 4  [^2] Additionally, routes are generated for slot 3-8 and 3-10 with shadow separately after modifying the sun azimuth angle and altitude angle. For example, when random seed 11 is assigned to slot 3-8 in scene generation, 6 routes are for training set and 2 routes are for validation set. The same for seed 27. 12 / 4
Slot 3-10 with shadow, seed 12, 28 12 / 4  [^2] Additionally, routes are generated for slot 3-8 and 3-10 with shadow separately after modifying the sun azimuth angle and altitude angle. For example, when random seed 11 is assigned to slot 3-8 in scene generation, 6 routes are for training set and 2 routes are for validation set. The same for seed 27. 12 / 4
Slot 3-8 with shadow, seed 91, 107, 123, 139 (3 / 1)  [^3] For slot 3-8 and 3-10 in Gen 2A, the routes with shadow in the slots are not generated separately, but are included in the cases with random seeds ranging from 16 to 143 mentioned above. Therefore, these routes are not counted in the total.
Slot 3-10 with shadow, seed 28, 44, 60, 76 (3 / 1)  [^3] For slot 3-8 and 3-10 in Gen 2A, the routes with shadow in the slots are not generated separately, but are included in the cases with random seeds ranging from 16 to 143 mentioned above. Therefore, these routes are not counted in the total. (1 / 1)  [^4] The routes for training and validation are assigned seeds 28 and 44, respectively, both of which are included in the cases mentioned above.
Total 96 / 32 192 / 64 216 / 72 192 / 64 248 / 104
[…]
5 Experiments
5.1 Experimental setup
The network architecture keeps the same as the original one and hyperparameters are slightly different. The network was trained on 4 NVIDIA RTX 3090 (total VRAM: 96 GB) for dataset Gen 1A, 1B, 1C and 2A, and on 4 NVIDIA RTX 4090 (total VRAM: 96 GB) for dataset Gen 2B, both with a batch size of 16. Adam optimizer is utilized with a weight decay of $1\times 10^{-4}$, while the beta values are 0.9 and 0.999 as default. The performance of models trained on 5 datasets with different epochs and initial learning rate will be compared.
5.1.1 Iteration of datasets
Gen 1A has 128 routes with random scenes, with 96 for training and 32 for validation. Each parking slot has only 1 random seed to generate a scene. Random seeds 0 to 15 are allocated to 16 parking slots in order. For each parking slot 8 routes are collected, where 6 routes are designated for the training set, and 2 routes are designated for the validation set. The initial positions of the 6 training routes are distributed as far-left, middle-left, near-left, near-right, middle-right, and far-right from a bird’s-eye view. For the 2 validation routes, their initial positions are relatively random, with the constraint that one is positioned on the left and the other on the right.
Gen 1B inherited 128 trajectories from Gen 1A and introduced an additional 128 trajectories, with random seeds 16–31 used for scene generation. The settings of initial positions and the split of training and validation set remain the same.
Building upon Gen 1B, Gen 1C further incorporated 16 trajectories for slot 3-8 with shadow and 16 trajectories for slot 3-10 with shadow. 24 of them are for training and 8 of them are for validation. The initial position settings and the division of the training and validation sets remain unchanged.
Similar to Gen 1A, 128 trajectories were generated for Gen 2A using random seeds 0 to 15. The difference is that the straight-line distance before reversing was shortened to prevent the vehicle from driving out of the parking lot. To enhance the diversity, more random seeds are used to generate scenes for the extra 128 routes, as Table 3 illustrates.
Table 3: Random seed and initial position assignment w.r.t weather settings in training and validation sets in Gen 2A
Dataset Weather 0 Weather 1 Weather 2
Training set 6 initial positions: 0 to 15 far-left: 16 to 31 near-right: 80 to 95
middle-left: 32 to 47 middle-right: 96 to 111
near-left: 48 to 63 far-right: 112 to 127
Validation set flexible: 0 to 15 flexible: 64 to 79 flexible: 128 to 143
[…]
00footnotetext: Each random seed is assigned to one of 16 parking slots with different weather settings to ensure a diverse dataset. Weather 0 is default weather, in which the shadow of streetlight is projected on slot 3-9. In Weather 1, the sun angles are set to project the shadow on slot 3-10, while in Weather 2, the shadow is on slot 3-8.
Compared to Gen 2A, Gen 2B eliminates the selection of random seeds ranging from 48 to 143. It retains 160 routes from Gen 2A, comprising 128 routes associated with seeds 0 to 15, 16 routes with seeds 16 to 31, and 16 routes with seeds 32 to 47. Including these inherited routes, there are 128 routes assigned to seeds 16 to 31, while 64 routes are assigned to seeds 32 to 47. For these 64 routes, each slot is allocated 2 routes for the training set and 2 routes for the validation set. The initial positions of the 2 training routes are set to far-left and far-right, whereas those of the 2 validation routes are positioned at middle-left and middle-right. Furthermore, 32 additional routes are generated for slots 3-8 and 3-10 with shadows, following the same initialization distribution and random seed selection strategy as the corresponding part in Gen 1C.
5.2 Metrics
The metrics used for model testing remain the same as the original ones.
Target Success Rate (TSR) represents the probability of the ego vehicle successfully parking in the designated slot. A parking attempt is considered successful if the vehicle’s center is within 0.6 meters horizontally and 1 meter longitudinally from the slot’s center, with an orientation deviation of no more than 10 degrees.
Target Failure Rate (TFR) refers to the probability of the ego vehicle reaching the target parking slot but with errors exceeding acceptable limits.
Non-Target Rate (NTR) represents the probability that the ego vehicle parks in a slot other than the designated target. NTR is the sum of the Non-Target Success Rate (NTSR) and Non-Target Failure Rate (NTFR), both of which depend on whether the error in parking into a non-target slot is acceptable.
Collision Rate (CR) refers to the probability of a collision occurring during the parking process.
Outbound Rate (OR) indicates the probability that the ego vehicle exits the parking area.
Timeout Rate (TR) refers to the probability that the ego vehicle fails to park successfully within a specified time limit. The original authors combine OR and TR and refers to their sum as TR.
Average Position Error (APE) is the mean error between the ego vehicle’s final position and the center of the target slot in successful parking instances.
Average Orientation Error (AOE) is the average difference between the ego vehicle’s final yaw angle and the desired orientation of the target slot in successful parking instances.
Average Parking Time (APT) represents the mean duration taken for successful parking maneuvers.
Average Inference Time (AIT) is the average time spent on network inference per step.
5.3 Results
To test the model efficiently, each model is required to accomplish 96 parking tasks (1 test epoch), with 6 tasks assigned to each of the 16 test parking slots, each task having a different initial position for the vehicle. The best model trained on Gen 2B will be tested for 4 test epochs with 384 tasks to ensure the stability of the test.
5.3.1 Comparison of Gen 1A, 1B, and 1C
The model trained on the Gen 1A dataset performed poorly in testing, achieving a TSR of only 44.792% after 155 epochs of training. For comparison, the model trained on the Gen 1B dataset achieved a TSR of 65.625% in testing after 155 epochs with the same learning rate of $1\times 10^{-4}$, benefiting from the doubled dataset size and increased scenario diversity.
Table 4: Gen 1A’s Performance
Metric Lr = $1\times 10^{-4}$
Epochs 120 155
val_loss 0.713 0.706
TSR(%) 40.625 44.792
[…]
A comparative experiment on different initial learning rates was conducted using the Gen 1B dataset. An initial learning rate of $5\times 10^{-4}$ caused oscillations in the training loss, so no testing was performed. Meanwhile, an initial learning rate of $5\times 10^{-5}$ results in a TSR of only around 50%, which indicates an appropriate learning rate should be chosen between these two values. The model trained with initial learning rate $Lr=7.5\times 10^{-5}$ after 195 epochs has the highest TSR of 71.875%.
Table 5: Comparison of GEN 1B’s Performance under Different Initial Learning Rates
Metric Lr = $1\times 10^{-4}$ Lr = $7.5\times 10^{-5}$ Lr = $5\times 10^{-5}$
Epochs val_loss TSR(%) Epochs val_loss TSR(%) Epochs val_loss TSR(%)
Early 115 0.737 57.292
Mid 155 0.725 65.625 155 0.700 65.625 155 0.721 50.000
Late 195 0.695 71.875 200 0.720 51.042
[…]
As Table 6 demonstrates, with the additional parking trajectories in shadowed slots, Gen 1C achieved a higher TSR in shadowed parking slots and outperformed Gen 1B in overall TSR. Specifically, with an initial learning rate of $1\times 10^{-4}$, the model trained on Gen 1C for 150 epochs achieved a TSR 7.25 percentage points higher than the model trained on Gen 1B for 155 epochs. Similarly, with an initial learning rate of $7.5\times 10^{-5}$, the model trained on Gen 1C for 155 epochs outperformed the Gen 1B model trained for the same number of epochs by 11.46 percentage points.
In Gen 1C with an initial learning rate of $1\times 10^{-4}$, the model made no progress in TSR between 150 and 190 training epochs. In contrast, with an initial learning rate of $7.5\times 10^{-5}$, the TSR improved by 3.13 percentage points between training from 65 to 155 epochs, but there is no improvement after that.
Table 6: Comparison of GEN 1C’s Performance under Different Initial Learning Rates
Metric Lr = $1\times 10^{-4}$ Lr = $7.5\times 10^{-5}$
Epochs val_loss TSR(%) Epochs val_loss TSR(%)
Early 65 0.799 73.958
Mid 150 0.752 72.917 155 0.764 77.083
Late 190 0.746 72.917 200 0.757 75.000
[…]
5.3.2 Comparison of Gen 2A and 2B
Based on the experience gained from Gen 1, an initial learning rate of $7.5\times 10^{-5}$ will be chosen for Gen 2, with a maximum of 155 epochs of training, to achieve a balance between model performance and training duration. The result is demonstrated in Table 7.
In Gen 2A, increasing the number of additional scenes without increasing the dataset size did not have a positive effect on model performance, as the TSR of 63.542% was 13.54 percentage points lower than the best result achieved by Gen 1C. The scene composition of Gen 2B has 16 more scenes compared to Gen 1C, ensuring that the additional scenes have sufficient parking route data. The model based on Gen 2B achieves the same TSR (77.083%) as that of Gen 1C. Due to the shorter straight-line distance before braking in Gen 2, the model performs better on parking slots near the edge of the parking lot. In Gen 1C, the TSR for slots 2-1 and 3-1 is only 16.667%, whereas in Gen 2B, the TSR for these slots is 66.667% and 50%, respectively.
Table 7: Comparison of Gen 2A and Gen 2B’s Performance
Metric Gen 2A Gen 2B
Epochs 150 130
val_loss 0.9550 0.8089
TSR(%) 63.542 77.083
[…]
00footnotetext: The initial learning rate is set as $7.5\times 10^{-5}$ during training on both datasets.
5.3.3 Details of Gen 2B testing
The preliminary test result of model on Gen 2B is illustrated in Table 8. The experiments consist of 96 tests in 16 different scenes. By observing the real-time video of the model’s inference and control of the vehicle, it is found that the vehicle’s speed is too high, leading to unintended movement towards non-target parking slots or even collisions. To address this issue, the speed for both gear 1 and reverse (R) was limited to 7.5 km/h. Additionally, a closed-loop experiment consisting of 384 test cases was conducted to evaluate the effectiveness of this adjustment. The result in Table 9 demonstrates that the model trained based on Gen 2B demonstrates robust performance, achieving a TSR of 85.156%. Compared to the original speed limit set by the author, the stricter speed limit resulted in an increase of 1.417 seconds in APT. However, this adjustment led to a 6.69% reduction in APE and a 14.76% decrease in AOE. More importantly, CR decreased from 9.375% to 2.865%, improving safety.
Table 8: Closed-loop results including 96 test cases for model trained on Gen 2B, with the speed limit set as default.
TaskIdx TSR TFR NTSR NTFR CR OR TR APE AOE APT AIT
2-1 66.667 0.000 0.000 0.000 0.000 33.333 0.000 0.277 0.821 20.458 0.078
2-3 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.278 0.209 19.844 0.078
2-5 83.333 0.000 0.000 0.000 16.667 0.000 0.000 0.273 0.315 20.673 0.078
2-7 66.667 16.667 0.000 0.000 16.667 0.000 0.000 0.259 0.159 19.383 0.078
2-9 50.000 0.000 0.000 0.000 50.000 0.000 0.000 0.242 0.313 18.711 0.078
2-11 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.307 0.444 20.189 0.079
2-13 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.299 0.611 19.417 0.078
2-15 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.214 0.435 20.228 0.078
3-1 50.000 0.000 0.000 0.000 0.000 33.333 16.667 0.213 1.218 23.489 0.078
3-3 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.263 0.236 19.528 0.078
3-5 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.286 0.180 20.711 0.079
3-7 66.667 16.667 0.000 0.000 0.000 0.000 16.667 0.207 0.256 20.708 0.087
3-9 33.333 0.000 0.000 0.000 66.667 0.000 0.000 0.209 0.382 19.950 0.090
3-11 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.279 0.289 21.094 0.084
3-13 66.667 0.000 0.000 0.000 0.000 0.000 33.333 0.242 0.239 22.900 0.079
3-15 50.000 0.000 0.000 0.000 0.000 0.000 50.000 0.225 0.175 23.322 0.078
Avg 77.083 2.083 0.000 0.000 9.375 4.167 7.292 0.254 0.393 20.663 0.080
[…]
Table 9: Closed-loop results including 384 test cases for model trained on Gen 2B, with the speed limit set as 7.5 km/h.
TaskIdx TSR TFR NTSR NTFR CR OR TR APE AOE APT AIT
2-1 54.167 0.000 8.333 0.000 0.000 20.833 16.667 0.259 0.586 22.645 0.079
2-3 91.667 0.000 0.000 0.000 0.000 0.000 8.333 0.258 0.193 22.283 0.079
2-5 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.253 0.182 21.131 0.079
2-7 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.257 0.141 21.233 0.079
2-9 70.833 0.000 4.167 0.000 20.833 0.000 4.167 0.250 0.207 21.748 0.079
2-11 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.239 0.450 21.347 0.081
2-13 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.309 0.360 20.854 0.080
2-15 95.833 0.000 0.000 0.000 0.000 0.000 4.167 0.262 0.481 22.539 0.079
3-1 62.500 0.000 0.000 0.000 0.000 37.500 0.000 0.245 0.868 21.551 0.079
3-3 95.833 0.000 0.000 0.000 4.167 0.000 0.000 0.168 0.324 21.498 0.079
3-5 91.667 0.000 0.000 0.000 0.000 0.000 8.333 0.306 0.505 22.469 0.080
3-7 95.833 0.000 0.000 0.000 0.000 0.000 4.167 0.190 0.134 21.637 0.080
3-9 75.000 0.000 0.000 0.000 20.833 0.000 4.167 0.237 0.232 21.693 0.080
3-11 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.196 0.171 22.703 0.079
3-13 62.500 0.000 0.000 0.000 0.000 0.000 37.500 0.122 0.255 24.681 0.079
3-15 66.667 0.000 0.000 0.000 0.000 0.000 33.333 0.242 0.266 23.268 0.079
Avg 85.156 0.000 0.781 0.000 2.865 3.646 7.552 0.237 0.335 22.080 0.079
[…]
5.3.4 Parking performance compared with baseline and original expert
According to Table 10, the model trained on the new dataset Gen 2B can essentially reproduce the performance of the baseline. Although TSR is 6.25 percentage points lower and APT increases by 6.36 seconds, APE decreases by 20%, and AOE decreases by 60.9%. Meanwhile, the new expert shows a significant improvement in accuracy compared to the original expert. Although APT increases by 1.99 seconds, APE decreases by 43.5% and AOE decreases by 90.0%.
Table 10: Performance comparison with baseline and original expert
TaskIdx TSR (%) TFR (%) CR (%) APE (m) AOE (deg) APT (s)
Baseline 91.41 2.08 2.08 0.30 0.87 15.72
Model trained on Gen 2B 85.16 0.00 2.87 0.24 0.34 22.08
Original Expert 100.00 0.00 0.00 0.23 0.48 14.96
New Expert 100.00 0.00 0.00 0.13 0.05 16.95
[…]
6 Conclusion
In this work we created an open-source dataset for E2E Parking and achieved a Target Success Rate comparable with the original authors. Moreover, the model trained on the dataset shows an enhanced parking precision with lower Average Position Error and Average Orientation Error. The iterative process of dataset construction provides insights for relevant researchers on how to optimize the dataset and improve training and testing results through methods such as parking operation strategies, scene selection, learning rate adjustment, and speed limitations.
Despite the promising results, this work has several limitations. First, the dataset was collected exclusively under optimal lighting conditions (sunny days at noon), which limits its ability to generalize to scenarios with different times of day or adverse weather conditions. Second, all obstacles present in the dataset are static, and therefore the model has not been exposed to dynamic objects, which may hinder its capacity for motion prediction and collision avoidance in real-world environments. Third, the current experiments are conducted in a simulated environment, which inevitably introduces a domain gap; the proposed algorithm may not transfer seamlessly to real-world applications due to sensor noise, actuation delays, and other environmental uncertainties.
Nonetheless, we believe this work provides a solid foundation for future research in end-to-end autonomous parking. By expanding the dataset to include more diverse environmental conditions and dynamic agents, refining the model architecture, and adapting and validating the proposed algorithm on real vehicles, future work can further advance the deployment of robust autonomous parking systems.